OpenStudy (anonymous):

Why does 1-.99999...=0

OpenStudy (bahrom7893):

it doesnt

OpenStudy (bahrom7893):

it doesnt

OpenStudy (saifoo.khan):

Disagree.

OpenStudy (anonymous):

then what does it equal

OpenStudy (bahrom7893):

0.0000.....1

OpenStudy (anonymous):

how can there be an ending 1

OpenStudy (bahrom7893):

depends on the number of 9s, but it doesn't equal 0, obviously u can round it, but it's never actually 0

OpenStudy (anonymous):

because 1/3=.333333 1/3+1/3+1/3=3/3 3/3=1 so .999999=1

OpenStudy (bahrom7893):

stephen is this just a joke?

OpenStudy (anonymous):

then what is it

OpenStudy (bahrom7893):

It is something very close to 0, but not 0

OpenStudy (anonymous):

so in a way... .9999999=1 and 1-1=0

OpenStudy (saifoo.khan):

2 - 2 = 0

OpenStudy (anonymous):

thanks madscientist

OpenStudy (anonymous):

You can use a proof to prove it... but that doesn't mean it's right... math has problems.

OpenStudy (anonymous):

and np ^.^

OpenStudy (anonymous):

oh it is zero all right

OpenStudy (anonymous):

HOW CAN THAT BE!!